OpenStudy (darkbluechocobo):

Help with confirming true or false question about graphing tangent functions

OpenStudy (darkbluechocobo):

The question is. The vertical asymptotes of the tangent function are located at the same values of x for which the cosine function is equal to 0. I believe it is true because of the following statement: Because\[\tan \theta=\frac{ \sin \theta }{ \cos \theta }\]anywhere the value of θ causes cos θ to be 0, the value of tangent is undefined. In fact, these are values where the graph of tangent will have vertical asymptotes.

OpenStudy (solomonzelman):

I don't really get the statement, but yes, any value of \(\large\color{black}{ \displaystyle \theta }\) that makes \(\large\color{black}{ \displaystyle \cos\theta=0 }\) is the vertical asympotote.

OpenStudy (darkbluechocobo):

Well I believe the statement was saying if Cos Theta=0 and that the tangent would have vertical asymptotes :p

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

anywhere where \(\large\color{black}{ \displaystyle \cos \theta }\) is equal to zero, that is a vertical asymptote.

OpenStudy (darkbluechocobo):

Thank you :D